OpenStudy (anonymous):

Describe the symmetry, if any, of the graph of the function x f(x) = ____ x^2 + 5 . A) no symmetry B) origin symmetry C) x-axis symmetry D) y-axis symmtery Eliminate

OpenStudy (michele_laino):

is your function like this: \(y=x^2+5\)

OpenStudy (mathmale):

Yes, clarity is an issue here. Describe the symmetry, if any, of the graph of the function f(x) = x+x^2 + 5 ??? Is that the function you're dealilng with?

OpenStudy (anonymous):

i edit for u @Michele_Laino

OpenStudy (michele_laino):

hint: if I make this transformation: \(x \rightarrow -x\), then the function \(f(x)\) will transform according this rule: \[\Large f\left( { - x} \right) = \frac{{ - x}}{{{{\left( { - x} \right)}^2} + 5}} = - \frac{x}{{{x^2} + 5}} = - f\left( x \right)\] so, what can you conclude?

OpenStudy (michele_laino):

hint: when I can write such formula, namely \(f(-x)=-f(x)\), then I say that \(f(x)\) is an \(odd \) function for example \(f(x)=x^3\) is an odd function, and its graph is like this: 

OpenStudy (anonymous):

i dont know i use math way

OpenStudy (michele_laino):

as we can see the function \(f(x)=x^3\), is symmetric with respect to the origin, since for any point \((x,y)\) I have its symmetric with respect to the origin:  so, what is the right option?